OpenStudy (anonymous):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (15, 45) and (150, 60), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points? (Write your answer in simplest form, using / for a fraction bar if needed.)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Do you remember the slope formula?

jimthompson5910 (jim_thompson5910):

If not, then it's m = (y2-y1)/(x2-x1) Tell me what you get

OpenStudy (anonymous):

mmmmm i get 0.111111111.......

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = ( 60- 45)/(150-15) m = 15/135 m = 1/9 So the slope is 1/9 which in decimal form is 0.1111

jimthompson5910 (jim_thompson5910):

so you are correct, but they want the fraction

OpenStudy (anonymous):

ohh ok lol so i put 1/9

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

Thank you! (;

jimthompson5910 (jim_thompson5910):

anytime